### v2

# Roberto Bruno, Roberto De Prisco, and Ugo Vaccaro

Abstract
In this paper, we consider the problem of constructing optimal average-length binary codes under the constraint that each codeword must contain at most $D$ ones, where $D$ is a given input parameter. We provide an $O(n^{2}D)$-time complexity algorithm for the construction of such codes, where $n$ is the number of codewords. We also describe several scenarios where the need to design these kinds of codes naturally arises. Our algorithms allow us to construct both optimal average-length prefix binary codes and optimal average-length alphabetic binary codes. In the former case, our $O(n^{2}D)$-time algorithm substantially improves on the previously known $O(n^{2+D})$-time complexity algorithm for the same problem. We also provide a Kraft-like inequality for the existence of (optimal) variable-length binary codes, subject to the above-described constraint on the number of 1’s in each codeword.
Index Terms—Variable-Length Codes, Alphabetic Codes, Binary Search, Dynamic Programming, Kraft-inequality.
I Introduction
Variable-length codes find plenty of applications in data compression [41]; additionally, variable-length codes are deeply interconnected with search theory. Specifically, any search procedure that identifies objects within a given space by sequentially performing appropriate tests inherently defines a variable-length encoding for the elements in that space. To illustrate, each possible test outcome can be represented using a distinct symbol from a finite alphabet. The concatenation of these encoded test outcomes, which ends in identifying an object $x$, provides a valid encoding of $x$, for any $x$ in the search space. The relationships between variable-length codes and search procedures have been thoroughly investigated; the main results are presented in the excellent monographies [1, 2, 35].
Among search procedures, the binary search stands out as an important algorithmic technique to solve the computational problem whereby an unknown element $x$ is located within a finite search space $S$, and the objective is to identify $x$ by performing queries that yield Yes/No answers. These kind of queries typically fall into one of two categories: comparison queries or membership queries. In the case of comparison queries, the unknown element $x$ resides in an ordered set that can be represented as $[n]=\{1,\ldots,n\}$ The queries then take the form: ”Is the unknown number $x$ greater than $j$, where $j\in[n]$?” (see [30] for an encyclopedic treatment of the subject). For membership queries, the search space $S$ may be unordered. The queries are phrased as: ”Does the unknown number $x$ belong to a subset $A\subseteq S$?” This latter problem has also been extensively studied (see, for example, [1, 2, 33]) and has significant connections to a variety of fields, including Exact Learning [10], Combinatorics [26], and Group Testing [13], to name a few. It is well known that binary search procedures that proceed with comparison queries are mathematically equivalent to binary alphabetic codes, and that binary search procedures that proceed with membership queries are equivalent to binary prefix codes (see, e.g., [1]).
Among the numerous versions of the basic binary search problem, we are interested in the following variation that arose in the context of efficient leader election in asynchronous anonymous networks [31]. We first describe the problem, then we show how it gives rise to our problem of constructing alphabetic binary codes with the constraint that each codeword must contain at most a given number of $D$ ones. We quote verbatim from [31], apart from the names of the parameters
“Consider the case when the unknown number is a positive integer in the interval $[1,n]$, and the total amounts $D$ of admissible overestimates is predetermined. The $\langle n,*,D\rangle$– game consists of determining the unknown number by asking as few questions as possible”.
The questions are of the type: “Is the unknown number greater than $j\in[1,n]$?”. With the term overestimate, the authors of [31] intend the situation in which the search algorithm, receives a ‘No’ answer to the query. In fact, if this is the case, it means that the value $j$ is greater than the value of the unknown number one seeks, hence, $j$ is an overestimate of the sought number. The authors of [31] also pointed out that their $\langle n,*,D\rangle$–game (as above defined) corresponds to the bounded-searching problem with variable-cost comparisons studied by Bentley and Brown in [5] and remarked that their results about $\langle n,*,D\rangle$–games lead to improved solutions for the algorithmic problems studied in [5].
In different words, the $\langle n,*,D\rangle$–search game problem consists in determining the best worst-case comparison-based search algorithm $\cal A$ that determines the unknown element in $[1,n]$, under the constraints that $\cal A$ never performs a sequence of comparison with more than $D$ overestimates of the unknown number. Using the well-known correspondence between (deterministic) comparison-based search procedures and binary trees [33], one can restate the $\langle n,*,D\rangle$–search game problem of [31] as the problem of determining the minimum possible height of any $n$-leaf binary tree, satisfying the constraints:
  1. 1.
    any root-to-leaf path contains at most $D$ right-pointing edges (assuming the right-pointing edges correspond to the No answers),
  2. 2.
    the leaves of the tree, read from the leftmost to the rightmost leaf, are in one-to-one correspondence with the elements of the ordered set $[n]=\{1,\ldots,n\}$.
Additionally, using the equally well-known correspondence between binary trees and variable-length binary codes (and assuming to encode each No answers with 1, and each each Yes answers with 0), one can see the $\langle n,*,D\rangle$–search game problem of [31] as the problem of determining an alphabetic encoding (because of 2. above) of the elements in $[n]$ such that: a) the encoding has minimum maximum codeword length, b) the encoding satisfies the constraint that each codeword must contain at most $D$ ones. In the rest of the paper, we will use interchangeably both the languages of codes and of trees, since they refer to the same objects, essentially.
A strictly related problem was considered by Dolev et al. in [12]. More precisely, the authors of [12] considered the problem of determining the minimum possible average height of any $n$-leaf binary tree that satisfies the following constraints:
  1. 1’.
    any root-to-leaf path contains at most $D$ right–pointing edges,
  2. 2’.
    the leaves of the tree, (not necessarily read from the leftmost to the rightmost leaf), are in one-to-one correspondence with the elements of the set $[n]$,
  3. 3’.
    the search space $[n]$ is endowed with a probability distribution $p=(p_{1},\ldots,p_{n})$, where $p_{i}$ is the probability that the unknown element one wants to determine is equal to $i\in[n].$
The authors of [12] were motivated by the goal of quantifying the time–energy trade-off in message transmissions between mobile hosts and mobile support stations. In the same paper [12] the authors give a $O(n^{2+D})$-time complexity algorithm to construct a minimum average height tree that satisfies the required properties 1’, 2’, and 3’. By the known correspondence between binary trees and search algorithms that operate by executing membership queries (see, e.g. [2], Ch. 1.9), one can see that any minimum-average height binary tree, satisfying 1’, 2’, and 3’, gives an average-case optimal search strategy in which: a) tests correspond to arbitrary tests of the type: “Does the unknown number belong to a subset $A\subseteq[n]$?” and b) the number of tests that can receive a No answer is upper bounded by $D$. It should be clear, at this point, that the problem can be equivalently formulated as that of constructing a minimum average length prefix code, in which each codeword contains at most $D$ ones.
We remark that problems of the kind described above arise in many other different scenarios (e.g., [34, 37, 38, 39, 40, 42, 43, 44, 45, 49]) where different test outcomes might have different costs. In our case, a Yes outcome costs 0, while a No has some cost $>0$ (like in the classic egg-dropping puzzle [28], for example), and one seeks average-case optimal algorithms (with respect to the number of performed tests) with a fixed bound on the worst-case cost of the algorithm. Related problems are also studied in [20, 25] and references therein quoted.
I-A Our results
In Section III we consider the scenario of the $\langle n,*,D\rangle$–search games studied in [31], where now the search space $[n]$ is endowed with a probability distribution $p=(p_{1},\ldots,p_{n})$, and $p_{i}$ is the probability that the unknown element is equal to $i\in[n].$ We design an $O(n^{2}D)$ time algorithm to find the best average case comparison-based search algorithm $\cal A$ that is capable of determining the unknown element in $[n]$, under the constraints that $\cal A$ never performs a sequence of comparisons with more than $D$ overestimates of the unknown number. Equivalently, our algorithm constructs a minimum average height binary tree satisfying properties 1. and 2. above, for the input probability distribution $p=(p_{1},\ldots,p_{n})$. Our main technical tool is the celebrated Knuth-Yao Dynamic Programming–speedup through quadrangle–inequality [29, 48].
Successively, in Section IV we derive the following result. Let $\{\ell_{1},\ldots,\ell_{n}\}$ be a multiset of positive integers, and $D\geq 1$ be an integer (by the definition, $\{\ell_{1},\ldots,\ell_{n}\}$ is unordered). We give a necessary and sufficient condition for the existence of an $n$-leaf full binary tree $T$ (i.e., in which any internal node has two children) such that:
  1. a)
    the multiset of all root-to-leaf path lengths is equal to $\{\ell_{1},\ldots,\ell_{n}\}$,
  2. b)
    any root-to-leaf path of $T$ contains at most $D$ right–pointing edges.
The interest in the above result is twofold. First, it represents a natural extension of the classical Kraft inequality (see, eg., Thm. 5.1 of [1]) that establishes a necessary and sufficient condition for the existence of an $n$-leaf full binary tree with multiset of root-to-leaf path lengths equal to $\{\ell_{1},\ldots,\ell_{n}\}$. Secondly, the above result will be instrumental to the derivation of an improvement of the result by Dolev et al. [12] that we have discussed before. More precisely, in Section V we design a novel algorithm to construct a minimum average height tree that satisfies properties 1’, 2’, and 3’; our algorithm has $O(n^{2}D)$ time complexity, considerably improving on the $O(n^{2+D})$ algorithm presented in [12]. Note that the parameter $D$ can possibly assume any value in the interval $[1,\log n]$; therefore, our algorithm is the first polynomial in $n$ time algorithm to solve the problem discussed before.
We conclude this section by mentioning that our work can also be seen as an extension of well-known classical results, in the sense that our first problem corresponds to the study of binary alphabetic codes [23] in which all codewords have a bounded number $D$ of 1’s, and the second problem corresponds to the extension of the Huffman encoding [22] to the same constrained scenario just mentioned. As such, our work fits into the long line of research devoted to the construction of minimum average variable-length codes satisfying given structural properties (see, e.g., [15, 18, 19, 47]).
II Preliminaries
We find it convenient to exploit the known equivalence between comparison-based search procedures, and binary alphabetic codes (see [1, 2]).
Let $S=\{s_{1},\ldots,s_{n}\}=[n]$ be the set of symbols, ordered according to the standard order. A binary alphabetic code is a mapping $w:[n]\longmapsto\{0,1\}^{+}$, enjoying the following two properties:
  1. 1.
    the mapping $w:[n]\longmapsto\{0,1\}^{+}$ is order-preserving, where the order relation on the set of all binary strings $\{0,1\}^{+}$ is the alphabetical order,
  2. 2.
    no codeword $w(s)$ is prefix of another $w(s^{\prime})$, for any $s,s^{\prime}\in[n]$, $s\neq s^{\prime}$.
In the following sections, we will occasionally see an alphabetic code as a binary tree $T$ where each left-pointing edge is labeled by bit 0 and each right-pointing edge is labeled by bit 1. Each leaf of the tree corresponds to a different symbol $s\in S$ and the order-preserving property of alphabetic codes implies that the leaves of $T$, read from the leftmost leaf to the rightmost one, appear in the order $1<2<\ldots<n$.
Given a probability distribution $p=(p_{1},\dots,p_{n})$, where $p_{i}$ is the probability of symbol $s_{i}$, for $i=1,\ldots,n$, we denote with the symbol $\mathbb{E}[C]$ the average length of an alphabetic code $C=\{w(s):s\in S\}$ for $S$, defined as $\mathbb{E}[C]=\sum_{i=1}^{n}p_{i}\ell_{i},$ where $\ell_{i}$ is the length of the codeword $w(s_{i})$.
In Section I we interpreted the $\langle n,*,D\rangle$–search games of [31] in terms of constrained $n$-leaf binary trees. That said, we can translate our algorithmic problems into the language of alphabetic codes as follows. For any binary string $w\in\{0,1\}^{+}$, let us denote with ${\lvert w\rvert_{1}}$ the number of 1’s that appear in $w$. The problem of designing the best average case comparison-based search algorithm $\cal A$ to determine the unknown element in $[n]$, under the constraints that $\cal A$ never performs a sequence of comparison with more than $D$ overestimates of the unknown number, is equivalent to solving the following optimization problem:
$\min_{C\mbox{ \footnotesize alphabetic}}\mathbb{E}[C]=\min_{C\mbox{ \footnotesize alphabetic}}\sum_{i=1}^{n}p_{i}\,\ell_{i},$ (1)
[…]
where $\ell_{i}$ is the length of $w(s_{i})$. In other words, one seeks an alphabetic code of minimum average length in which each codeword contains no more than $D$ ones.
III A Dynamic Programming approach
In this section, we first provide a straightforward Dynamic Programming algorithm to compute a solution to the optimization problem (1). Subsequently, we employ the Knuth-Yao Dynamic Programming–speedup to get a better algorithm.
Let $C(i,j,w)$ denote the minimum average length of any alphabetic code (tree) for the symbols $s_{i},\ldots,s_{j}$, with $1\leq i\leq j\leq n$, in which each codeword contains at most $w$ ones. We say that $C(i,j,w)$ is the cost of the optimal alphabetic tree. The solution of the optimization problem (1) corresponds to the problem of determining the value $C(1,n,D)$. Let $\text{c}(i,j)=p_{i}+\cdots+p_{j}$ denote the sum of the probabilities associated with the symbols $s_{i},\ldots,s_{j}$. We express the cost $C(i,j,w)$ as a function of the cost of its subtrees. By considering all the possible ways of splitting the sequence of symbols $s_{i},\ldots,s_{j}$ into a left and a right subtree (with at least a node in each subtree), we get the following recurrence relation:
$C(i,j,w)=\begin{cases}0&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ if }i=j,\\ \infty&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ if }i<j\text{ and }w=0,\\ \text{c}(i,j)+\min_{i<k\leq j}\{C(i,k-1,w)+\\ \qquad C(k,j,w-1)\}&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ otherwise.}\end{cases}$ (2)
We remind that the right subtree is reached from a right-pointing edge of the root; therefore, all the corresponding codewords one obtains begin with symbol 1, and this accounts for the value $w-1$ in the parameters of the right subtree cost.
From the expression of (2), one can easily see that $C(1,n,D)$, can be computed in $O(n^{3}D)$ time and $O(n^{3}D)$ space. We explicitly write the Algorithm 1 that does that, since it will be useful for discussing the improvement in the time complexity, presented in the next Section III-A. To compute also the optimal alphabetic code (and not only its cost) the algorithm keeps track of the indices on which the optimal partition occurs. This is done with the values $R(i,j,w)$’s, i.e., $R(i,j,w)$ corresponds to the index $k$ for which $C(i,j,w)=\text{c}(i,j)+C(i,k-1,w)+C(k,j,w-1).$
Input: Symbols $S=\{s_{1},\dots,s_{n}\}$, $p=(p_{1},\dots,p_{n})$ the associated probability distribution and an integer value $D$
1
2for $i\leftarrow 1$ to $n$ do
3       for $j\leftarrow i+1$ to $n$ do
4             $C(i,j,0)=\infty$
5
6
7for $w\leftarrow 0$ to $D$ do
8       for $i\leftarrow 1$ to $n$ do
9             $C(i,i,w)=0$
10            $R(i,i,w)=i$
11
12
13for $w\leftarrow 1$ to $D$ do
14       for $i\leftarrow 1$ to $n-1$ do
15             $C(i,i+1,w)=p_{i}+p_{i+1}$
16            $R(i,i+1,w)=i$
17
18for $w\leftarrow 1$ to $D$ do
19       for $s\leftarrow 2$ to $n-1$ do
20             for $i\leftarrow 1$ to $n-s$ do
21
                  $j=i+s$ ;
                    // Proceed by diagonals
22
23                  $min=\infty$
24                  $mindex=-1$
25                  for $k\leftarrow i+1$ to $j$ do
26                         if $C(i,k,w)+C(k+1,j,w-1)\leq min$ then
27                               $min=C(i,k-1,w)+C(k,j,w-1)$
28                              $mindex=k$
29
30                  $C(i,j,w)=\left(\sum_{k=i}^{j}p_{k}\right)+min$
31                  $R(i,j,w)=mindex$
32
33
Output: $C(1,n,D)$ and $R(\cdot,\cdot,\cdot)$
Algorithm 1
III-A An improved algorithm
The time complexity of Algorithm 1 can be improved by a factor of $n$ by using a technique introduced by Knuth in [29] and generalized by Yao [48]. For such a purpose, let us recall some preliminary definitions and results.
Let $R_{i,j,w}$ be the largest index where the minimum is achieved in the definition of $C(i,j,w)$. We recall that $R_{i,j,w}$ corresponds exactly to the value $R(i,j,w)$ computed in Algorithm 1. Moreover, let us use $C_{k}(i,j,w)$ to denote the cost of an optimal alphabetic code under the constraint that the first splitting of symbols $s_{i},\ldots,s_{j}$ is performed into $s_{i},\dots,s_{k-1}$ and $s_{k},\ldots,s_{j}$, i.e, $C_{k}(i,j,w)$ satisfies
$C_{k}(i,j,w)=\text{c}(i,j)+C(i,k-1,w)+C(k,j,w-1).$ (3)
We recall that a two-argument function $f(\cdot,\cdot)$ satisfies the Quadrangle Inequality [48] if it holds that $f(i,j)+f(i^{\prime},j^{\prime})\leq f(i^{\prime},j)+f(i,j^{\prime}),\ \forall 1\leq i\leq i^{\prime}\leq j\leq j^{\prime}\leq n.$ A function $f(\cdot,\cdot)$ is said to be monotone if it holds that $f(i,j^{\prime})\geq f(i^{\prime},j),\quad\forall 1\leq i\leq i^{\prime}\leq j\leq j^{\prime}\leq n.$ One can easily see that $\text{c}(i,j)=p_{i}+\cdots+p_{j}$ is both monotone and satisfies the Quadrangle Inequality (actually, with the equality sign).
We also recall the work of Borchers and Gupta [6] in which they showed that the Quadrangle Inequality holds for a large class of functions. Let $F(i,j,r)$ be the minimum cost of solving a problem on inputs $i,i+1,\dots,j$ with “resource” $r$ (in our case, the resource will be the maximum number of 1’s one can have in each codeword). Borchers and Gupta showed that the Quadrangle Inequality holds for all functions $F(i,j,r)$ defined as follows:
$F(i,j,r)=\begin{cases}w(i),&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ if }i=j\text{ and }r\geq 0,\\ \min_{i\leq k<j}\{aF(i,k,f(r))+bF(k+1,j,g(r))\\ \qquad+h(i,k,j)\},&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ if }f(r)\geq 0\text{ and }g(r)\geq 0,\\ \mbox{undefined},&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\text{ otherwise,}\end{cases}$ (4)
where $w(i)$ denotes the cost of $F(i,i,r)$, $f(\cdot)$ and $g(\cdot)$ are two functions such that $r\geq f(r)$ and $r\geq g(r)$, i.e., they are assuming that the resource does not increase, $a$ and $b$ are two positive integers, and the function $h(i,k,j)$ is the cost of dividing the problem $F(i,j,r)$ at the splitting point $k$. Moreover, the function $h$ must satisfy the following conditions that are similar to the quadrangle inequality: for $i\leq j\leq t<s\leq l$ and $i\leq s<l$, if $t\leq k$ it holds that
$h(i,t,s)-h(j,t,s)+h(j,k,l)-h(i,k,l)\leq 0$ (5)
and
$h(j,k,l)-h(i,k,l)\leq 0.$ (6)
While if $k<t$, it holds that
$h(j,t,l)-h(j,t,s)+h(i,k,s)-h(i,k,l)\leq 0$ (7)
and
$h(i,k,s)-h(i,k,l)\leq 0.$ (8)
The authors of [6] proved the following result.
Lemma 1 ([6]).
For each $r\geq 0$, the function $F$ defined in (4) satisfies the Quadrangle Inequality, that is, $\forall 1\leq i\leq i^{\prime}\leq j\leq j^{\prime}\leq n$ it holds that
$F(i,j,r)+F(i^{\prime},j^{\prime},r)\leq F(i,j^{\prime},r)+F(i^{\prime},j,r).$ (9)
Using Lemma 1 we can prove the following result.
Corollary 1.
For each $0\leq w\leq D$, our function $C$ defined in (2) satisfies the Quadrangle Inequality, that is, $\forall 1\leq i\leq i^{\prime}\leq j\leq j^{\prime}\leq n$, it holds that
$C(i,j,w)+C(i^{\prime},j^{\prime},w)\leq C(i,j^{\prime},w)+C(i^{\prime},j,w).$ (10)
Proof.
We need to show that the function $C$ is a special case of (4). For such a purpose, let $w(i)=0$ for each $i=1,\dots,n$, and let $a=b=1$. While as functions $f$ and $g$ we take $f(x)=x$ and $g(x)=x-1$. Finally, we define the function $h(i,k,j)=\text{c}(i,j)$ for each $i\leq k<j$, where $\text{c}(i,j)=p_{i}+\dots+p_{j}$. One can see that the function $h$ defined in this way satisfies the conditions (5-8). In fact, for $i\leq j\leq t<s\leq l$ and $i\leq s$, if $t\leq k$ it holds that
$h(i,t,s)$ $-h(j,t,s)+h(j,k,l)-h(i,k,l)$ (11)
$=\text{c}(i,s)-\text{c}(j,s)+\text{c}(j,l)-\text{c}(i,l)=0$ (12)
and
$h(j,k,l)-h(i,k,l)=\text{c}(j,l)-\text{c}(i,l)<0.$ (13)
Similarly, if $k<t$, it holds that
$h(j,t,l)$ $-h(j,t,s)+h(i,k,s)-h(i,k,l)$ (14)
$=\text{c}(j,l)-\text{c}(j,s)+\text{c}(i,s)-\text{c}(i,l)=0$ (15)
and
$h(i,k,s)-h(i,k,l)=\text{c}(i,s)-\text{c}(i,l)<0.$ (16)
Since we have shown that the function $C$ is a special case of (4), from Lemma 1 we get that the function $C$ satisfies the Quadrangle Inequality. ∎
By using Corollary 1 we can show the following result which is a generalization of the crucial property called “monotonicity of the roots” in [29]. This result will be the key step for the improved algorithm.
Theorem 1.
For each $0\leq w\leq D$ it holds that
$\forall 1\leq i\leq j\leq n\quad R_{i,j,w}\leq R_{i,j+1,w}\leq R_{i+1,j+1,w}.$ (17)
Proof.
Let $1\leq w\leq D$ (for $w=0$ the theorem is trivially true). The claim is trivially true also when $i=j$. Let us assume that $i<j$. We will show that $R_{i,j,w}\leq R_{i,j+1,w}$, the argument for $R_{i,j+1,w}\leq R_{i+1,j+1,w}$ follows by symmetry. Since $R_{i,j,w}$ is the largest index where the minimum is achieved in the definition of $C(i,j,w)$, it is sufficient to show that $\forall i<k\leq k^{\prime}\leq j$ it holds that
$C_{k^{\prime}}(i,j,w)\leq C_{k}(i,j,w){\implies}C_{k^{\prime}}(i,j+1,w)\leq C_{k}(i,j+1,w).$ (18)
We will prove that (18) holds by showing the following inequality, for all $i<k\leq k^{\prime}\leq j,$
$C_{k}(i,j,w)-C_{k^{\prime}}(i,j,w)\leq C_{k}(i,j+1,w)-C_{k^{\prime}}(i,j+1,w).$ (19)
In fact, let us assume that (19) is true. Then, whenever $C_{k^{\prime}}(i,j,w)\leq C_{k}(i,j,w)$ of (18) holds, from (19) we get that
$0\leq C_{k}(i,j,w)-C_{k^{\prime}}(i,j,w)\leq C_{k}(i,j+1,w)-C_{k^{\prime}}(i,j+1,w).$ (20)
From (20), since $0\leq C_{k}(i,j+1,w)-C_{k^{\prime}}(i,j+1,w)$, we obtain that $C_{k^{\prime}}(i,j+1,w)\leq C_{k}(i,j+1,w)$, that is, (18) holds too. To prove (19) let us rewrite it as follows
$C_{k}(i,j,w)+C_{k^{\prime}}(i,j+1,w)\leq C_{k}(i,j+1,w)+C_{k^{\prime}}(i,j,w).$ (21)
By expanding all four terms of (21) we get (cfr., see (3))
c $(i,j)+C(i,k-1,w)+C(k,j,w-1)+\text{c}(i,j+1)+C(i,k^{\prime}-1,w)+C(k^{\prime},j+1,w-1)$
$\leq\text{c}(i,j+1)+C(i,k-1,w)+C(k,j+1,w-1)+\text{c}(i,j)+C(i,k^{\prime}-1,w)+C(k^{\prime},j,w-1).$
Grouping the terms of the above inequality, we obtain
$C(k,j,w-1)+C(k^{\prime},j+1,w-1)\leq C(k,j+1,w-1)+C(k^{\prime},j,w-1).$ (22)
This is the quadrangle inequality for the function $C$ with parameter $w-1$, with $k\leq k^{\prime}\leq j<j+1$, and from Corollary 1 we know that it holds true. Thus, this concludes the proof. ∎
Due to Theorem 1, instead of searching from $i+1$ to $j$ (line 18 of Algorithm 1), it is sufficient to search from $R(i,j-1,w)$ through $R(i+1,j,w)$. Therefore line 18 of Algorithm 1 can be substituted by “for $k\leftarrow R(i,j-1,w)$ to $R(i+1,j,w)$”. Exactly as argued by Knuth (see the remark after the Corollary, p. 19 of [29]), when $C(i,j,w)$ is being calculated with this modification, only $R(i+1,j,w)-R(i,j-1,w)+1$ cases need to be examined. Summing for fixed $j-i$ gives a telescoping series showing that the total amount of work of each iteration of the loop at line 12 of the modified Algorithm 1 is proportional to $O(n^{2})$, at worst. This decreases the total time complexity of the algorithm from $O(n^{3}D)$ to $O(n^{2}D)$.
IV A Kraft-like condition for the existence of constrained binary trees
Let $\{\ell_{1},\ldots,\ell_{n}\}$ be a multiset of positive integers, and $D$ a positive integer. In this section, we derive a necessary and sufficient condition for the existence of an $n$-leaf full binary tree $T$ such that: a) the (unordered) multiset of all root-to-leaf path lengths is equal to $\{\ell_{1},\ldots,\ell_{n}\}$, and b) any root-to-leaf path of $T$ contains at most $D$ right–pointing edges. In addition to its intrinsic interest (indeed, the result represents a natural extension of the classical Kraft inequality (Thm. 5.1 of [1])), it will be essential for the derivation of an improvement of the result by Dolev et al. [12] that we have anticipated in Section I-A. Using the natural correspondence between binary trees and binary prefix codes, we will first derive a sufficient condition for the existence of binary prefix codes with the constraint that the number of 1’s in each codeword is bounded by a given parameter $D$. We point out that there are many extensions of the Kraft inequality to situations where the codes have to satisfy additional constraints (besides that of being prefix), see [3, 9, 11, 17, 24, 27, 36, 46]. However, none of the papers quoted above deals with the extension we derive in this section.
Lemma 2.
Let $\{\ell_{1},\ldots,\ell_{n}\}$ be a multiset of positive integers, $D$ a positive integer and $L=\max_{i}\ell_{i}$. Let $N_{j}$ count the number of integers in $\{\ell_{1},\ldots,\ell_{n}\}$ that are equal to $j$, for $j=1,\dots,L$. Then, there exists a binary prefix code with codeword lengths $\ell_{1},\ldots,\ell_{n}$, such that each codeword does not contain more than $D$ $1$’s, if the following inequalities hold:
$N_{j}\leq\sum_{i=0}^{D-1}\binom{j}{i}+\binom{j-1}{D-1}-M_{j+1},\quad\forall j=1,\ldots,L,\ \mbox{ where }$ (23)
$M_{j+1}=\begin{cases}0\quad&\mbox{ if }j\geq L,\\ \left\lceil\frac{\displaystyle N_{j+1}+M_{j+2}}{\displaystyle 2}\right\rceil&\mbox{ if }1\leq j<L.\end{cases}$ (24)
Proof.
As per established convention, we assume that $\binom{j}{i}=0$ if $i>j$. Let $T$ be the binary tree of maximum depth $L$ such that: 1) $T$ is full, 2) $T$ has the maximum number of nodes, and 3) any root-to-leaf path of $T$ has at most $D$ right-pointing edges
We label each left-pointing edge of $T$ with 00 and each right-pointing edge with $1$. The set of all binary words one gets by reading the sequences of binary labels from the root of $T$ to each leaf gives a binary prefix code $P$ in which ${\lvert w\rvert_{1}}\leq D$, for each $w\in P.$ We now show that the number of nodes at the level $j$ of such a tree $T$ is exactly equal to
$\sum_{i=0}^{D-1}\binom{j}{i}+\binom{j-1}{D-1}\quad\forall j=1,\ldots,L.$ (25)
First of all, we recall that each path in the tree can be seen as a binary string. Therefore, (25) holds because the first term $\sum_{i=0}^{D-1}\binom{j}{i}$ counts the number of binary strings of length $j$, each containing at most $D-1$ ones. The number of binary strings of length $j$ containing exactly $D$ ones that one obtains, by reading the root-to-leaves path in $T$ is equal to $\binom{j-1}{D-1}$. This is true because any binary string of length $j$ that one obtains by reading the root-to-leaves path in $T$ and that contains $D$ ones must end with 1. In fact, let us consider a binary string of length $j$ that contains exactly $D$ ones and does not end with 1. In the path from the root to the node associated with such a binary string, there will be at least one internal node that has only exactly one child. Therefore, such a node cannot belong at level $j$ of the binary tree $T$ since, by hypothesis, $T$ is full.
We show that for an arbitrary list of positive integers $\ell_{1},\ldots,\ell_{n},D$ that satisfy the inequalities (23), there is a prefix code $C$ with $\ell_{1},\ldots,\ell_{n}$ as codeword lengths, obeying the constraint that ${\lvert w\rvert_{1}}\leq D$, for each $w\in C.$ The idea of the proof is to build the code by taking as codewords the binary sequences one gets by following the paths from the root to nodes of $T$ of levels $\ell_{1},\ldots,\ell_{n}$, starting from level $L=\max_{i}\ell_{i}$ (i.e., by constructing the longest codewords first). From (23) we know that
$N_{L}\leq\sum_{i=0}^{D-1}\binom{L}{i}+\binom{L-1}{D-1}.$ (26)
Hence, there are enough available nodes in the level $L$ of $T$ to construct $N_{L}$ codewords of length $L$. In particular, it will be convenient (for future purposes) to construct such codewords by choosing the first $N_{L}$ nodes, in sequence, starting from the left-most one. The constraint that the obtained codewords $w$’s have to satisfy (i.e., that ${\lvert w\rvert_{1}}\leq D$) is automatically satisfied since in the tree $T$ any root-to-leaf path has at most $D$ right-pointing edges, each labeled with $1$.
The nodes at level $L-1$ that are parents of the chosen nodes at level $L$ cannot be used anymore in the construction of the code, to comply with the prefix constraint of $C$. Therefore, we need to ignore such nodes at level $L-1$ and their number is exactly $M_{L}=\lceil N_{L}/2\rceil.$ Thus, the number of nodes at level $L-1$ that can be still used is
$\sum_{i=0}^{D-1}\binom{L-1}{i}+\binom{L-2}{D-1}-M_{L}.$ (27)
From (23) and (24) we know that
$N_{L-1}\leq\sum_{i=0}^{D-1}\binom{L-1}{i}+\binom{L-2}{D-1}-M_{L},$ (28)
therefore we can iterate the same process to construct the desired number of codewords of length $L-1$. We choose the first $N_{L-1}$ available nodes at level $L-1$ (that are all the nodes at level $L-1$ that give rises to not-prefixes of the chosen nodes at level $L$) in sequence starting from the left-most one. Again, we have to ignore all the nodes at level $L-2$ that are ancestors of the chosen nodes at level $L$ and $L-1$. This number is exactly
$M_{L-1}=\left\lceil\frac{\displaystyle N_{L-1}+M_{L}}{\displaystyle 2}\right\rceil.$ (29)
From (23) and (24) we know that
$N_{j}\leq\sum_{i=0}^{D-1}\binom{j}{i}+\binom{j-1}{D-1}-M_{j+1},\quad\forall j=1,\ldots,L.$ (30)
We can iterate this operation for each level of $T$. Thus, we can construct a prefix code with codeword lengths $\ell_{1},\ldots,\ell_{n}$ in which each codeword does not contain more than $D$ ones. ∎
To clarify the procedure for the construction of the prefix code described in the previous Lemma, let us consider a concrete example.
Example 1.
Let $\langle 4,2,3,4,3,3,3\rangle$ be the list of lengths associated to the symbols $s_{1},\ldots,s_{7}$ and $D=2$. By the definitions, we get the values: $N_{4}=2,N_{3}=4,N_{2}=1$ and $N_{1}=0$. Let us first of all see that for such values the inequalities (23) hold. Indeed:
$N_{4}=2$ $\leq\sum_{i=0}^{1}\binom{4}{i}+\binom{3}{1}=8,$ (31)
$N_{3}=4$ $\leq\sum_{i=0}^{1}\binom{3}{i}+\binom{2}{1}-\left\lceil\frac{2}{2}\right\rceil=5,$ (32)
$N_{2}=1$ $\leq\sum_{i=0}^{1}\binom{2}{i}+\binom{1}{1}-\left\lceil\frac{4+1}{2}\right\rceil=1,$ (33)
$N_{1}=0$ $\leq\sum_{i=0}^{1}\binom{1}{i}+\binom{0}{1}-\left\lceil\frac{1+3}{2}\right\rceil=0.$ (34)
Since the inequalities hold, we can construct the prefix code with the required list of lengths $\langle 4,2,3,4,3,3,3\rangle$. For such a purpose, let us start from the complete binary tree $T$ of depth $4$ in which, for any node $v$, the path from the root to the node $v$ does not contain more than $D=2$ edge with label $1$ (Figure 1).
Figure 1: The complete binary tree $T$ of depth $4$
Since $N_{4}=2$ we need two nodes at the level $4$ of $T$ to construct the codewords for the symbols $s_{1}$ and $s_{4}$, respectively. As described in Lemma 2, we take the left-most ones and we get the tree in Figure 2. Note that is the inequality (31) that assures us that we can choose such nodes at the level $4$.
$s_{1}$$s_{4}$
Figure 2: The tree $T$ after choosing the nodes for the codewords of length 4
Now, since $N_{3}=4$ we need to take four nodes at level $3$ for the construction of the codewords for the symbols $s_{3},s_{5},s_{6}$ and $s_{7}$. However, we have to ignore the nodes that are prefixes of the chosen codewords of length $4$. Note that we can do this since the values $N_{i}$’s satisfy the required inequalities.
$s_{1}$$s_{4}$$s_{3}$$s_{5}$$s_{6}$$s_{7}$
Figure 3: The tree after choosing the nodes for the codewords of length $4$ and $3$
Since $N_{2}=1$ there is only one codeword of length 2 that we need to choose. Therefore, we take the left-most node at level $2$ that is still “available” as shown in Figure 4.
$s_{1}$$s_{4}$$s_{3}$$s_{5}$$s_{6}$$s_{7}$$s_{2}$
Figure 4: The tree after choosing the nodes for the codewords of length $4,3$ and $2$
Finally, since $N_{1}=0$ we do not need to take any node at level $1$. So, the tree that we get in the end is shown in Figure 5.
$s_{1}$$s_{4}$$s_{3}$$s_{5}$$s_{6}$$s_{7}$$s_{2}$
Figure 5: The final tree after the choice of all codewords
The prefix code $w:\{s_{1},\ldots,s_{7}\}\to\{0,1\}^{+}$ that we get from the tree in Figure 5 is the following: $w(s_{1})=0000$, $w(s_{2})=11$, $w(s_{3})=001$, $w(s_{4})=0001$, $w(s_{5})=010$, $w(s_{6})=011$ and $w(s_{7})=100$.
$\Box$
The condition presented in Lemma 2 is only a sufficient condition for the existence of binary prefix codes with a given multiset of codeword-lengths, satisfying the constraint that the number of ones in each codeword is bounded by the parameter $D$. If we consider multisets containing codeword-lengths of prefix codes whose tree representation corresponds to a full binary tree (i.e., for which it holds that $\sum_{i=1}^{n}2^{-\ell_{i}}=1$ or, equivalently, any internal node has exactly two children [1]), then the condition of Lemma 2 becomes necessary too, as shown in the following Lemma.
Lemma 3.
Let $\ell_{1},\ldots,\ell_{n}$ be the codeword lengths of a prefix code $C$ in which each codeword does not contain more than $D$ ones and whose tree representation corresponds to a full binary tree, then the following inequalities hold:
$N_{j}\leq\sum_{i=0}^{D-1}\binom{j}{i}+\binom{j-1}{D-1}-M_{j+1},\quad\forall j=1,\ldots,L,$ (35)
where $M_{j}$ and $N_{j}$ are defined as in Lemma 2 and $L=\max_{i}\ell_{i}$.
Proof.
First of all, let us recall that
$\sum_{i=0}^{D-1}\binom{j}{i}+\binom{j-1}{D-1}$ (36)
is the maximum number of nodes that a complete binary tree (in which any root-to-leaf path has at most $D$ right-pointing edges) can have at level $j=1,\dots,L$. Moreover, since we are considering only full binary trees, one can see that $M_{j}$, for each $j=2,\dots,L$, corresponds to the minimum number of nodes at level $j-1$ that are necessarily prefixes of some nodes associated with codewords of length $\geq j$, regardless of how such nodes are chosen. Putting all together, since by hypothesis, $\ell_{1},\dots,\ell_{n}$ are the codeword lengths of a prefix code $C$ in which each codeword does not contain more than $D$ ones and whose binary tree representation is full, we have that the number of leaves at level $j$ in the binary tree of $C$, which is $N_{j}$, cannot be greater than
$\sum_{i=0}^{D-1}\binom{j}{i}+\binom{j-1}{D-1}-M_{j+1},$ (37)
for each $j=1,\dots,L$. This holds because (37) is obtained by subtracting from the total number of nodes at level $j$ (counted by (36)) the minimum number of nodes at level $j-1$ that are necessarily prefixes of some nodes associated with codewords of length $\geq j$ (counted by $M_{j+1}$). In other words, (37) represents the maximum number of leaves at level $j=1,\dots,L$, that a complete binary tree — where any root-to-leaf path has at most $D$ right-pointing edges — has still “available” (i.e., they are not prefixes of the nodes associated with codewords of length $\geq j$). ∎
V Optimal Average-Length Constrained Trees
In this section, we turn our attention to the problem considered by Dolev et al. in [12]; we recall it here for the reader’s convenience. The problem consists in determining the minimum possible average height of any $n$-leaf binary tree that satisfies the following constraints: a) any root-to-leaf path contains at most $D$ right–pointing edges, b) the leaves of the tree, (not necessarily read from the leftmost to the rightmost leaf), are in one-to-one correspondence with the elements of the set $[n]$, c) the set $[n]$ is endowed with a probability distribution $p=(p_{1},\ldots,p_{n})$. By using the known equivalence between binary prefix codes and binary trees, the problem is equivalent to the following optimization problem:
$\min_{C\mbox{ \footnotesize prefix}}\mathbb{E}[C]$ $=\min_{C\mbox{ \footnotesize prefix}}\sum_{i=1}^{n}p_{i}\,\ell_{i},$ (38)
[…]
where $\ell_{i}$ is the length of $w(s_{i})$. We shall prove that the problem (38) can be solved in $O(n^{2}D)$ time, improving on the $O(n^{2+D})$ time algorithm given in [12].
The essence of our proof consists in showing that, under the special condition that the probability distribution $p$ is ordered in a non-decreasing fashion, then the value of an optimal solution to the problem (38) is equal to the value of an optimal solution to (1).
Theorem 2.
Let $S=\{s_{1},\ldots,s_{n}\}$ be a set of symbols, $p=(p_{1},\dots,p_{n})$ be a probability distribution on $S$ such that $p_{1}\leq\ldots\leq p_{n}$, and let $D$ be a strictly positive integer. Then, there exists an optimal prefix code $C$ for $S$ and $p$ such that: 1) each codeword of $C$ does not contain more than $D$ 1’s, 2) $\ell_{1}\geq\ldots\geq\ell_{n}$, where $\ell_{i}$ is the length of $w(s_{i})$, and 3) in the tree representation of $C$ the leaves associated to the symbols $s_{1},\ldots,s_{n}$ appear consecutively, from left to right.
Proof.
The existence of an optimal (i.e., with minimum average length) prefix code $C^{\prime}$ in which the codeword lengths are ordered in opposite fashion with respect to the probabilities ordering is obvious. Therefore, properties 1) and 2) are true. Moreover, since $C^{\prime}$ is an optimal code, its tree representation is a full binary tree. Therefore, from Lemma 3 we have that the codeword lengths $\ell_{1},\ldots,\ell_{n}$ of $C^{\prime}$ satisfy the inequalities
$N_{j}\leq\sum_{i=0}^{D-1}\binom{j}{i}+\binom{j-1}{D-1}-M_{j+1},\quad\forall j=1,\ldots,\ell_{1}.$ (39)
On the other hand, from Lemma 2 we know that there exists a prefix code $C$ for $S$ with codewords lengths $\ell_{1},\ldots,\ell_{n}$ such that in the tree representation of $C$ the leaves associated to the symbols $s_{1},\ldots,s_{n}$ appear consecutively from left to right (for more details, please see the argument right after (26) in the proof of Lemma 2). More explicitly, in the constructive proof of Lemma 2 we build the code tree $T$ by always taking the leftmost available nodes, starting from the lowest level suitable node, and proceeding iteratively according to the decreasing order of the $\ell_{1},\ldots,\ell_{n}$. This concludes the proof. ∎
The crucial consequence of Theorem 2 is that one can proceed as follows to solve the optimization problem (38). First, one observes that the value of a solution to (38) does not depend on the componentwise ordering of the probability distribution $p=(p_{1},\dots,p_{n})$ (this is analogous to the classical Huffman problem to find a minimum average length prefix code, for example). Therefore, we can assume that $p_{1}\leq\ldots\leq p_{n}$. Under this assumption, Theorem 2 asserts that there exists a binary tree $T$ representing an optimal solution to (38) in which the symbols $s_{1},\ldots,s_{n}$ (that is, ordered according to decreasing probabilities) appear on the leaves of $T$ consecutively, from left to right. Said equivalently, $T$ represents an optimal alphabetic code for the (now ordered) set of symbols $\{s_{1},\ldots,s_{n}\}$. We have shown in Section III-A that we can find an optimal alphabetic code in time $O(n^{2}D)$, for any arbitrary (but otherwise fixed) ordering on the set of symbols $\{s_{1},\ldots,s_{n}\}$. All together, this shows that an optimal solution to the problem (38) can be found in time $O(n^{2}D)$ by first ordering the probability distribution and then running the algorithm presented Section III-A.
V-A When the constraint is really a constraint?
The constraint that at most $D$ ones must appear in each codeword might prevent the construction of the (otherwise unrestricted) optimal tree. However, if $D$ is large enough, the best tree satisfying the constraint can be the same as the optimal (unrestricted) one. For the case of alphabetic trees establishing the smallest value of $D$ that does not prevent the construction of the optimal tree is straightforward. Indeed, it is easy to identify probability distributions (for example, a decreasing dyadic distribution) for which any optimal alphabetic code must have a codeword with $n-1$ ones. Hence, any value of $D$ strictly less than $n-1$ is a constraint that prevents the construction of a tree with the same cost of an optimal unrestricted one.
For the case of prefix codes, establishing a similar result is not immediate. In the next theorem we prove that the smallest value of $D$ that does not prevent the construction of a tree whose cost is as the cost of an optimal unrestricted tree is $D=\log_{2}n$.
Theorem 3.
Let $S=\{s_{1},\ldots,s_{n}\}$ be a set of symbols and $p=(p_{1},\dots,p_{n})$ a probability distribution on $S$ with $p_{1}\leq\dots\leq p_{n}$. There exists an optimal prefix tree $T^{*}$ for $S$, in which each codeword does not contain more than $\log_{2}n$ ones.
Proof.
We prove the theorem by induction.
Base case. For $n\leq 3$ one can see that the theorem holds, as shown in Figure 6.
$s_{1}$
(a) Optimal prefix tree for $n=1$
$s_{1}$ $s_{2}$
(b) Optimal prefix tree for $n=2$
$s_{1}$$s_{2}$$s_{3}$
(c) Optimal prefix tree for $n=3$
Figure 6:
Inductive step. Let $T^{*}$ be an optimal prefix tree for $S$ with $\ell_{1}\geq\ldots\geq\ell_{n}$, where
  • $\ell_{i}$ is the length of the codeword $w(s_{i})$ associated to symbol $s_{i}$,
  • the leaves of $T^{*}$, associated to the symbols $s_{1},\ldots,s_{n}$, appear consecutively, from left to right.
We note that such a tree surely exists. Indeed, when the source symbol probability distribution $p=(p_{1},\dots,p_{n})$ is ordered, an optimal alphabetic code for $S$ is also an optimal prefix code for $S$. Let $T_{1}$ and $T_{2}$ be the left and the right subtree of the root of $T^{*}$. By the inductive hypothesis, we have that each codeword of $T_{1}$ does not contain more than $\log_{2}|T_{1}|$ ones, where $|T_{1}|$ is the number of leaves of $T_{1}$. Similarly, each codeword of $T_{2}$ does not contain more than $\log_{2}|T_{2}|$ ones, where $|T_{2}|$ is the number of leaves of $T_{2}$. Moreover, since $T^{*}$ is full, and since $\ell_{1}\geq\dots\geq\ell_{n}$ and the leaves associated with the symbols $s_{1},\ldots,s_{n}$ appear consecutively, from left to right, one gets that $|T_{1}|\geq|T_{2}|$. In fact, let us assume by contradiction that $|T_{1}|<|T_{2}|$. Let $x$ be the depth of the highest leaf in $T_{1}$ and let $y$ be the depth of the lowest leaf in $T_{2}$. Since the trees are full, we have that $x\leq\log_{2}|T_{1}|$ and that $y\geq\log_{2}|T_{2}|$. From this, we get that
$y\geq\log_{2}|T_{2}|>\log_{2}|T_{1}|\geq x.$
However, $y>x$ leads to a contradiction since $\ell_{1}\geq\dots\geq\ell_{n}$.
Putting it all together, we get that each codeword of $T^{*}$ does not contain a number of 1’s greater than
$\max\{\log_{2}|T_{1}|,\log_{2}|T_{2}|+1\}$ $\leq\max\{\log_{2}|T_{1}|,\log_{2}\frac{n}{2}+1\}\quad\mbox{(since $|T_{2}|\leq|T_{1}|$)}$
$\leq\log_{2}n\quad\mbox{(since $|T_{1}|+|T_{2}|=n)$}$
This concludes the inductive step and thus the proof. ∎
VI Concluding Remarks
In this paper, we studied problems that can be seen as search problems in which different test outcomes (for the same test) can have different costs. Specifically, a ’No’ test outcome has a cost of 1, while a ’Yes’ test outcome has a cost of 0. We sought search strategies with the minimum average number of tests, under the constraint that the employed algorithm never incurs a total cost more than a given parameter $D$ for any single execution on input instances. We pointed out several scenarios where our problems naturally occur.
In this paper, we have provided efficient algorithms for the stated problems, extending and considerably improving on the known results in the literature. An interesting line of research is to develop our approach in such a way as to handle arbitrary variable test outcome costs (i.e., not necessarily 0/1). While the extension of the results of Section III is feasible, extending the results of Section V to the more general scenario is much more challenging. The source of difficulties seems to lie in the extension of the combinatorial analysis carried out in Section IV to the general variable test outcome costs. We leave this extension to future investigations.